OpenStudy (anonymous):

solve for d when 2x-2yd=y+xd that solved for y i need d

OpenStudy (anonymous):

2x-2yd=y+xd Arrange the variables alphabetically within the expression -2yd. This is the standard way of writing an expression. 2x-2dy=y+xd Reorder the polynomial 2x-2dy alphabetically from left to right, starting with the highest order term. -2dy+2x=y+xd Arrange the variables alphabetically within the expression xd. This is the standard way of writing an expression. -2dy+2x=y+dx Reorder the polynomial y+dx alphabetically from left to right, starting with the highest order term. -2dy+2x=dx+y Since y contains the variable to solve for, move it to the left-hand side of the equation by subtracting y from both sides. -2dy+2x-y=dx Reorder the polynomial -2dy+2x-y alphabetically from left to right, starting with the highest order term. -2dy-y+2x=dx Factor out the GCF of -y from the expression -2dy. -y(2d)-y=dx-2x Factor out the GCF of -y from the expression -y. -y(2d)-y(1)=dx-2x Factor out the GCF of -y from -2dy-y. -y(2d+1)=dx-2x Divide each term in the equation by (2d+1). -(y(2d+1))/(2d+1)=(dx)/(2d+1)-(2x)/(2d+1) Reduce the expression by canceling out the common factor of (2d+1) from the numerator and denominator. -(y<X>(2d+1)<x>)/(<X>(2d+1)<x>)=(dx)/(2d+1)-(2x)/(2d+1) Reduce the expression by canceling out the common factor of (2d+1) from the numerator and denominator. -y=(dx)/(2d+1)-(2x)/(2d+1) Combine all similar expressions in the polynomial. -y=(dx-2x)/(2d+1) Factor out the GCF of x from the expression dx. -y=(x(d)-2x)/(2d+1) Factor out the GCF of x from the expression -2x. -y=(x(d)+x(-2))/(2d+1) Factor out the GCF of x from dx-2x. -y=(x(d-2))/(2d+1) Multiply each term in the equation by -1. -y*-1=(x(d-2))/(2d+1)*-1 Multiply -y by -1 to get y. y=(x(d-2))/(2d+1)*-1 Multiply the rational expressions to get ((x*-1(d-2)))/((2d+1)). y=(x*-1(d-2))/(2d+1) Multiply x by -1 to get -x. y=-(x(d-2))/(2d+1)